OpenStudy (hulagirlmb):

Point: (7,4) Slope: 6 Can somebody help me write the equation?

OpenStudy (anonymous):

use this y-y1=m(x-x1) where m=6 x1=7 and y1=4

OpenStudy (hulagirlmb):

I don't get it

OpenStudy (anonymous):

its a formula that you have to use